# Solutions of the singular Yamabe problem near singular boundaries

### v1

@@ 2. Preliminaries
$D_{2}+D_{4}\leq C_{2}\frac{u_{V}r}{d^{2}}\left(1+A\beta u_{V}^{\beta-1}+Br-BC_{1}^{-\frac{4}{n-2}}C_{2}^{-1}n\right),$
@@
$A\beta u_{V}^{\beta-1}\leq\left\{\begin{aligned} AC_{1}^{\frac{2}{n-2}}d,&\quad n\geq 4,\\ 0,&\quad n=3,\end{aligned}\right$.
@@
$D_{2}+D_{4}$ $\leq\left\{\begin{aligned} C_{2}\frac{u_{V}r}{d^{2}}\left(1+AC_{1}^{\frac{2}{n-2}}r+Br-BC_{1}^{-\frac{4}{n-2}}C_{2}^{-1}n\right),&\quad n\geq 4,\\ C_{2}\frac{u_{V}r}{d^{2}}\left(1+Br-BC_{1}^{-\frac{4}{n-2}}C_{2}^{-1}n\right),&\quad n=3,\end{aligned}\right$.
$\leq C_{2}\frac{u_{V}r}{d^{2}}\left(1+\left((C_{3}C_{1}^{\frac{2}{n-2}}+1)r-C_{1}^{-\frac{4}{n-2}}C_{2}^{-1}n\right)B\right).$
@@
$(C_{3}C_{1}^{\frac{2}{n-2}}+1)r-C_{1}^{-\frac{4}{n-2}}C_{2}^{-1}n\leq-\frac{C_{1}^{-\frac{4}{n-2}}C_{2}^{-1}n}{2},\quad B\geq\frac{2C_{1}^{\frac{4}{n-2}}C_{2}}{n},$
@@
For any $s>0$ and $y\in B_{s}(0)$, we set $u_{s}(y)=(\frac{2s}{s^{2}-|y|^{2}})^{\frac{n-2}{2}}$. By directly calculations, we know
@@
for $s$ small enough. Therefore, choose $0<s<r_{2}$ small enough, we have that in $V\cap T(B_{R}(0))\cap B_{s}(0)$,
@@
where $f$ is Lipschitz and $f(0^{\prime})=0$. Set $F(\cdot)=\frac{n(n-2)}{4}(\cdot)^{\frac{n+2}{n-2}}$, and
$v^{\epsilon}(y)=v(y_{1},\cdots,y_{n-1},y_{n}+\epsilon),$
by (3.10) and $\mathcal{L}v=F(v)$ we have for any $y\in V\cap T(B_{R}(0))\cap B_{s}(0),$
@@
where $\varphi$ is between $w+2u_{s}$ and $v^{\epsilon}$. Recall $c=-\frac{n-2}{4(n-1)}S_{g}$, we may assume $|c|\leq C_{g}$ where $C_{g}>0$ is a constant depending only on $n$ and $g$. Besides, we can take $s>0$ small enough such that,
(3.11) $\frac{n(n+2)}{4}(w+2u_{s})^{\frac{4}{n-2}}>C_{g},\quad\mbox{ in }V\cap T(B_{R}(0))\cap B_{s}(0).$
Take $G=V\cap T(B_{R}(0))\cap B_{s}(0)\cap\{y:v^{\epsilon}(y)>(\frac{4C_{g}}{n(n+2)})^{\frac{n-2}{4}}\}$, by (3.11) we have
$F^{\prime}(\varphi)>C_{g},\quad\mbox{ in }G,$
which implies $F^{\prime}(\varphi)-c>0$ in $G$. Note $u_{s}=+\infty$ on $\partial B_{s}$, $w=+\infty$ on $\partial V\cap T(B_{R}(0))$, $0<s<<R$ and (3.11) tell us that $w+2u_{s}>v^{\epsilon}$ on $\partial G$, then, by
$a_{ij}(w+2u_{s}-v^{\epsilon})_{ij}+b_{i}(w+2u_{s}-v^{\epsilon})_{i}-(F^{\prime}(\varphi)-c)(w+2u_{s}-v^{\epsilon})\leq 0,\quad\mbox{ in }G,$
@@
$w+2u_{s}-v^{\epsilon}\geq 0,\quad\mbox{ in }G.$
On the other hand, by (3.11), in $V\cap T(B_{R}(0))\cap B_{s}(0)\cap\{y:v^{\epsilon}(y)\leq(\frac{4C_{g}}{n(n+2)})^{\frac{n-2}{4}}\}$ we have
$v^{\epsilon}\leq(\frac{4C_{g}}{n(n+2)})^{\frac{n-2}{4}}\leq w+2u_{s}.$
@@
$w+2u_{s}-v^{\epsilon}\geq 0,\quad\mbox{ in }V\cap T(B_{R}(0))\cap B_{s}(0),$
@@
(3.12) $v\leq w+2u_{s},\quad\mbox{ in }V\cap T(B_{R}(0))\cap B_{s}(0).$
Note in $B_{\frac{s}{2}}(0)$, $u_{s}\leq C(n,s)$, where $C(n,s)>0$ is a constant depending only on $n$ and $s$, then for any $0<s^{\prime}\leq\frac{s}{2},$ by (3.12) and (3.3) we have
@@
where $C>0$ is a constant depending only in $n$, size of exterior cones and $g$. By a process similar to that in the proof of [SW, Theorem 6.2] and note that we have (2.1)-(2.2) by Lemma 2.1, we can improve the estimate in dimension $3$ and get
$v-u_{V}\leq Cu_{V}|y|,\quad\mbox{ in }V\cap T(B_{R}(0))\cap B_{s^{\prime}}(0).$
@@
$v-u_{V}\geq-Cu_{V}|y|,\quad\mbox{ in }V\cap T(B_{R}(0))\cap B_{s^{\prime\prime}}(0),$
where $C>0$ is a constant depending only in $n$, size of exterior cones and $g$. Therefore, by Lemma 2.2,
$\left|\frac{u(x)}{u_{V}(Tx)}-1\right|\leq C|Tx|\leq Cd_{g}(x,0),\quad\mbox{ in }\Omega\cap B_{r_{0}}(0),$
@@
In this section, we study the first order expansion of local positive solutions of (1.1)-(1.2) near singular points on $\partial\Omega$ when $T$ has higher regularity. Formal calculations lead to an elliptic operator such as (A.2) or (A.8). However, we cannot use Theorem A.1 or Lemma A.2 directly since $c$ in (A.2) may be positive or $f$ in (A.9) may not belong to $L^{2}(\Sigma)$. Specifically, we study the following elliptic equation on $\Sigma$:
(4.1) $\rho^{2}\Delta_{\theta}\xi_{1}-\frac{n(n+2)}{4}\Big(1+\frac{n-4}{n+2}\rho^{2}\Big)\xi_{1}=\rho^{2}F,$
where $\rho$ is defined in (2.7) and $F\in C^{\infty}(\Sigma)$ satisfies
@@
Let $\Sigma\subset\mathbb{S}^{n-1}$ be a Lipschitz domain and $\xi\in C^{\infty}(\Sigma)$ be a positive function in $\Sigma$ satisfying (1.5)-(1.6). Then, for any $F\in C^{\infty}(\Sigma)$ with (4.2), there exists a solution $\xi_{1}$ of (4.1), and
(4.3) $|\xi_{1}|\leq C\xi\quad\text{on }\Sigma,$
where $C$ is a positive constant depending only on $n$, $\Sigma$, and $\bar{C}$. Moreover, if we further assume that $F$ is a linear combination of $\nabla_{\theta}\xi$ and $\nabla^{2}_{\theta}\xi$, then
(4.4) $d_{\Sigma}|\nabla_{\theta}\xi_{1}|+d^{2}_{\Sigma}|\nabla^{2}_{\theta}\xi_{1}|\leq C\xi\quad\text{on }\Sigma,$
where $C$ is a positive constant depending only on $n$, $\Sigma$, and $\bar{C}$.
@@
Set $\rho_{0}=\sqrt{n/2}$. Note that
(4.5) $-\frac{n(n+2)}{4}\Big(1+\frac{n-4}{n+2}\rho^{2}\Big)+\frac{(n-2)^{2}}{4}\rho^{2}+\frac{n(n-1)}{4}$
$\qquad=-\frac{3n}{4}+\rho^{2}<-\frac{3n}{4}+\frac{n}{2}=-\frac{n}{4}<0\quad\text{on }\{\rho<\rho_{0}\},$
@@
(4.6) $\begin{cases}\begin{aligned} c&=-\frac{n(n+2)}{4}\Big(1+\frac{n-4}{n+2}\rho^{2}\Big)\quad\text{on }\{\rho<\rho_{0}\},\\ c&\leq-\frac{(n-2)^{2}}{4}\rho^{2}-\frac{n(n-1)}{4}\quad\text{on }\Sigma.\end{aligned}\end{cases}$
Step 1. Set $L_{0}=\rho^{2}\Delta_{\theta}+c$, we solve
(4.7) $L_{0}\tilde{\xi}_{1}=\rho^{2}F.$
@@
Now, we can take $\psi=\xi$ and $\delta=n/4$ in Theorem A.1. Hence, we obtain a $\tilde{\xi}_{1}\in C^{\infty}(\Sigma)$ that satisfies (4.7) with
(4.8) $\Big\|\frac{\tilde{\xi}_{1}}{\xi}\Big\|_{L^{\infty}(\Sigma)}\leq\frac{1}{\delta}\Big\|\frac{\rho^{2}F}{\xi}\Big\|_{L^{\infty}(\Sigma)}\leq\frac{\bar{C}}{\delta},$
@@
Step 2. Set $L_{1}=\Delta_{\theta}-\frac{n(n+2)}{4\rho^{2}}$, we solve
(4.9) $L_{1}\bar{\xi}_{1}-\frac{n(n-4)}{4}\bar{\xi}_{1}=G,$
@@
$G=\tilde{\xi}_{1}\Big(\frac{c}{\rho^{2}}+\frac{n(n+2)}{4\rho^{2}}+\frac{n(n-4)}{4}\Big).$
By (4.6), we have $G=0$ on $\{\rho<\rho_{0}\}$ and $G\in C^{\infty}(\bar{\Sigma})$. Note that $n(n-4)/4\leq 0<\lambda_{1}$ for $n\geq 4$ and $n(n-4)/4=3/4<\lambda_{1}$ for $n=3$ by Lemma A.2(1) and Lemma A.2(3). Hence, by applying Lemma A.2(2) with $\lambda=-n(n-4)/4$ and $f=G$, we find a $\bar{\xi}_{1}\in C^{\infty}(\Sigma)$ satisfying (4.7) with
$\|\bar{\xi}_{1}\|_{H^{1}_{0}(\Sigma)}\leq C\|G\|_{L^{2}(\Sigma)}\leq C.$
@@
(4.10) $\|\bar{\xi}_{1}\|_{L^{\infty}(\Sigma)}\leq C(\|\bar{\xi}_{1}\|_{L^{2}(\Sigma)}+1)\leq C.$
Step 3. Take $\xi_{1}=\tilde{\xi}_{1}+\bar{\xi}_{1}$. By (4.7) and (4.9), we can verify that $\xi_{1}\in C^{\infty}(\Sigma)$ satisfies (4.1). In addition, by combining (4.8) and (4.10), we obtain (4.3). Now, we assume that $F$ is a linear combination of $\nabla_{\theta}\xi$ and $\nabla^{2}_{\theta}\xi$ and consider the estimates of $\nabla_{\theta}\xi_{1}$ and $\nabla^{2}_{\theta}\xi_{1}$. Recall that $\xi=\rho^{-(n-2)/2}\sim d_{\Sigma}^{-(n-2)/2}$. Note that for any $z\in B_{d_{\Sigma}(x)/2}(x)$,
@@
By applying the scaled $C^{1,\alpha}$-estimate to (1.5), we have, for any $x\in\Sigma$ and $\beta\in(0,1)$,
$d^{\beta}_{\Sigma}(x)[\xi]_{C^{\beta}(B_{d_{\Sigma}(x)/4}(x))}+d_{\Sigma}(x)\|\nabla_{\theta}\xi\|_{L^{\infty}(B_{d_{\Sigma}(x)/4}(x))}+d^{1+\beta}_{\Sigma}(x)[\nabla_{\theta}\xi]_{C^{\beta}(B_{d_{\Sigma}(x)/4}(x))}$
$\qquad\leq C(\|\xi\|_{L^{\infty}(B_{d_{\Sigma}(x)/2}(x))}+d^{2}_{\Sigma}(x)\|\xi^{\frac{n+2}{n-2}}\|_{L^{\infty}(B_{d_{\Sigma}(x)/2}(x))})$
$\qquad\leq C\|\xi\|_{L^{\infty}(B_{d_{\Sigma}(x)/2}(x))}.$
Next, by the scaled $C^{2,\alpha}$-estimate, we also obtain
$d^{2}_{\Sigma}(x)\|\nabla^{2}_{\theta}\xi\|_{L^{\infty}(B_{d_{\Sigma}(x)/8}(x))}+d^{2+\beta}_{\Sigma}(x)[\nabla^{2}_{\theta}\xi]_{C^{\beta}(B_{d_{\Sigma}(x)/8}(x))}$
$\qquad\leq C(\|\xi\|_{L^{\infty}(B_{d_{\Sigma}(x)/4}(x))}+d^{2}_{\Sigma}(x)\|\xi^{\frac{n+2}{n-2}}\|_{L^{\infty}(B_{d_{\Sigma}(x)/4}(x))}+d^{2+\beta}_{\Sigma}(x)[\xi^{\frac{n+2}{n-2}}]_{C^{\beta}(B_{d_{\Sigma}(x)/4}(x))})$
$\qquad\leq C(\|\xi\|_{L^{\infty}(B_{d_{\Sigma}(x)/4}(x))}+d^{2+\beta}_{\Sigma}(x)[\xi]_{C^{\beta}(B_{d_{\Sigma}(x)/4}(x))}\|\xi^{\frac{4}{n-2}}\|_{L^{\infty}(B_{d_{\Sigma}(x)/4}(x))})$
$\qquad\leq C\|\xi\|_{L^{\infty}(B_{d_{\Sigma}(x)/2}(x))}.$
Finally, by applying the scaled $C^{2,\alpha}$-estimate to (4.1) and using (4.2) and (4.3), we get
$d_{\Sigma}(x)\|\nabla_{\theta}\xi_{1}\|_{L^{\infty}(B_{d_{\Sigma}(x)/16}(x))}+d^{2}_{\Sigma}(x)\|\nabla^{2}_{\theta}\xi_{1}\|_{L^{\infty}(B_{d_{\Sigma}(x)/16}(x))}$
$\qquad\leq C(\|\xi_{1}\|_{L^{\infty}(B_{d_{\Sigma}(x)/8}(x))}+\frac{d^{2}_{\Sigma}(x)}{d^{2}_{\Sigma}(x)}\|\rho^{2}F\|_{L^{\infty}(B_{d_{\Sigma}(x)/8}(x))}+\frac{d^{2+\beta}_{\Sigma}(x)}{d^{2}_{\Sigma}(x)}[\rho^{2}F]_{C^{\beta}(B_{d_{\Sigma}(x)/8}(x))})$
@@
Let $\lambda_{1}$ be the first eigenvalue of $L_{1}$ in (A.8) and $\phi_{1}\in H^{1}_{0}(\Sigma)\cap C^{\infty}(\Sigma)\cap C(\bar{\Sigma})$ be the corresponding eigenfunction satisfying $\phi_{1}>0$ in $\Sigma$ with $\|\phi\|_{L^{2}(\Sigma)}=1$. Note that the existence of such a function $\phi_{1}$ is guaranteed by Lemma A.2.

### v2

@@ 2. Preliminaries
$D_{2}+D_{4}\leq C_{2}\frac{u_{V}r}{d^{2}}\left(1+A(\beta+1)u_{V}^{\beta-1}+Br-BC_{1}^{-\frac{4}{n-2}}C_{2}^{-1}n\right),$
@@
$A(\beta+1)u_{V}^{\beta-1}\leq\left\{\begin{aligned} 2AC_{1}^{\frac{2}{n-2}}d,&\quad n\geq 4,\\ AC_{1}d^{\frac{n-2}{2}},&\quad n=3,\end{aligned}\right$.
@@
$D_{2}+D_{4}$ $\leq\left\{\begin{aligned} C_{2}\frac{u_{V}r}{d^{2}}\left(1+2AC_{1}^{\frac{2}{n-2}}r+Br-BC_{1}^{-\frac{4}{n-2}}C_{2}^{-1}n\right),&\quad n\geq 4,\\ C_{2}\frac{u_{V}r}{d^{2}}\left(1+AC_{1}d^{\frac{n-2}{2}}+Br-BC_{1}^{-\frac{4}{n-2}}C_{2}^{-1}n\right),&\quad n=3,\end{aligned}\right$.
$\leq C_{2}\frac{u_{V}r}{d^{2}}\left(1+\left((2C_{3}C_{1}^{\frac{2}{n-2}}+1)r+C_{3}C_{1}r^{\frac{1}{2}}-C_{1}^{-\frac{4}{n-2}}C_{2}^{-1}n\right)B\right).$
@@
$(2C_{3}C_{1}^{\frac{2}{n-2}}+1)r_{2}+C_{3}C_{1}r_{2}^{\frac{1}{2}}-C_{1}^{-\frac{4}{n-2}}C_{2}^{-1}n\leq-\frac{C_{1}^{-\frac{4}{n-2}}C_{2}^{-1}n}{2},\quad B\geq\frac{2C_{1}^{\frac{4}{n-2}}C_{2}}{n},$
@@
For any $s>0$ and $y\in B_{s}(0)$, we set $u_{s}(y)=(\frac{2s}{s^{2}-|y|^{2}})^{\frac{n-2}{2}}$. By direct calculations, we know that
@@
for $s$ small enough. Therefore, by choosing $0<s<r_{2}$ small enough, we have that in $V\cap T(B_{R}(0))\cap B_{s}(0)$,
@@
where $f$ is Lipschitz and $f(0^{\prime})=0$. Set $F(\cdot)=\frac{n(n-2)}{4}(\cdot)^{\frac{n+2}{n-2}}$ and
$v^{\epsilon}(y)=v(y_{1},\cdots,y_{n-1},y_{n}+\epsilon).$
By (3.10) and $\mathcal{L}v=F(v)$, we have for any $y\in V\cap T(B_{R}(0))\cap B_{s}(0),$
@@
where $\varphi$ is between $w+2u_{s}$ and $v^{\epsilon}$. Recall $c=-\frac{n-2}{4(n-1)}S_{g}$, we may assume $|c|\leq C_{g}$, where $C_{g}>0$ is a constant depending only on $n$ and $g$. In addition, we can take $s>0$ small enough such that
(3.11) $\frac{n(n+2)}{4}(w+2u_{s})^{\frac{4}{n-2}}>C_{g}\quad\mbox{ in }V\cap T(B_{R}(0))\cap B_{s}(0).$
Set $G=V\cap T(B_{R}(0))\cap B_{s}(0)\cap\{y:v^{\epsilon}(y)>(\frac{4C_{g}}{n(n+2)})^{\frac{n-2}{4}}\}$, by (3.11), we have
$F^{\prime}(\varphi)>C_{g}\quad\mbox{ in }G,$
which implies $F^{\prime}(\varphi)-c>0$ in $G$. Note $u_{s}=+\infty$ on $\partial B_{s}$ and $w=+\infty$ on $\partial V\cap T(B_{R}(0))$. Since $0<s<<R$, it follows from (3.11) that $w+2u_{s}>v^{\epsilon}$ on $\partial G$. Then, by
$a_{ij}(w+2u_{s}-v^{\epsilon})_{ij}+b_{i}(w+2u_{s}-v^{\epsilon})_{i}-(F^{\prime}(\varphi)-c)(w+2u_{s}-v^{\epsilon})\leq 0\quad\mbox{ in }G,$
@@
$w+2u_{s}-v^{\epsilon}\geq 0\quad\mbox{ in }G.$
On the other hand, by (3.11), in $V\cap T(B_{R}(0))\cap B_{s}(0)\cap\{y:v^{\epsilon}(y)\leq(\frac{4C_{g}}{n(n+2)})^{\frac{n-2}{4}}\}$, we have
$v^{\epsilon}\leq\left(\frac{4C_{g}}{n(n+2)}\right)^{\frac{n-2}{4}}\leq w+2u_{s}.$
@@
$w+2u_{s}-v^{\epsilon}\geq 0\quad\mbox{ in }V\cap T(B_{R}(0))\cap B_{s}(0),$
@@
(3.12) $v\leq w+2u_{s}\quad\mbox{ in }V\cap T(B_{R}(0))\cap B_{s}(0).$
Note that in $B_{\frac{s}{2}}(0)$, $u_{s}\leq C(n,s)$, where $C(n,s)>0$ is a constant depending only on $n$ and $s$, then for any $0<s^{\prime}\leq\frac{s}{2},$ by (3.12) and (3.3) we have
@@
where $C>0$ is a constant depending only on $n$, the size of the exterior cones, and $g$. By a process similar to that in the proof of [20, Theorem 6.2] and note that we have (2.1)-(2.2) by Lemma 2.1, we can improve the estimate in dimension $3$ and get
$v-u_{V}\leq Cu_{V}|y|\quad\mbox{ in }V\cap T(B_{R}(0))\cap B_{s^{\prime}}(0).$
@@
$v-u_{V}\geq-Cu_{V}|y|\quad\mbox{ in }V\cap T(B_{R}(0))\cap B_{s^{\prime\prime}}(0),$
where $C>0$ is a constant depending only on $n$, the size of the exterior cones, and $g$. Therefore, by Lemma 2.2,
$\left|\frac{u(x)}{u_{V}(Tx)}-1\right|\leq C|Tx|\leq Cd_{g}(x,0)\quad\mbox{ in }\Omega\cap B_{r_{0}}(0),$
@@
In this section, we study the first order expansion of local positive solutions of (1.1)-(1.2) near singular points on $\partial\Omega$ when $T$ has higher regularity. Formal calculations lead to an elliptic operator such as (A.2) or (A.8). However, we cannot use Theorem A.1 or Lemma A.2 directly since $c$ in (A.2) may be positive or $f$ in (A.9) may not belong to $L^{2}(\Sigma)$. Specifically, for some constant $\sigma\in\mathbb{R}$ and positive integer $k$, we study the following elliptic equation on $\Sigma$:
(4.1) $\rho^{2}\Delta_{\theta}h-\frac{n(n+2)}{4}h+\sigma_{k}\rho^{2}h=\rho^{2}(F+\sigma\phi_{1}),$
where $\rho$ is defined in (2.7),
(4.2) $\sigma_{k}:=\Big(k-\frac{n}{2}+1\Big)\Big(k+\frac{n}{2}-1\Big),$
$\phi_{1}\in H^{1}_{0}(\Sigma)\cap C^{\infty}(\Sigma)\cap C(\bar{\Sigma})$ is the eigenfunction corresponding to the first eigenvalue $\lambda_{1}$ of $L_{1}$ in (A.8), satisfying $\phi_{1}>0$ on $\Sigma$ with $\|\phi_{1}\|_{L^{2}(\Sigma)}=1$, and $F\in C^{\infty}(\Sigma)$ satisfies
@@
Let $\Sigma\subsetneq\mathbb{S}^{n-1}$ be a Lipschitz domain. Suppose $\xi\in C^{\infty}(\Sigma)$ is a positive function on $\Sigma$ satisfying (1.5)-(1.6), and $F\in C^{\infty}(\Sigma)$ satisfies (4.3). For a given positive integer $k$, denote by $\sigma_{k}$ the constant defined in (4.2) with $\sigma_{k}\leq\lambda_{1}$.
$\mathrm{(i)}$ If $\sigma_{k}<\lambda_{1}$, then for $\sigma=0$, there exists a solution $h$ of (4.1);
$\mathrm{(ii)}$ If $\sigma_{k}=\lambda_{1}$, then for some $\sigma\in\mathbb{R}$, there exists a solution $h$ of (4.1).
In both cases,
(4.4) $|h|\leq C\xi\quad\text{on }\Sigma,$
where $C$ is a positive constant depending only on $n$, $k$, $\Sigma$, and $\bar{C}$. Moreover, if we further assume that for some constant $\beta\in(0,1)$,
(4.5) $d^{\beta}_{\Sigma}(x)[\rho^{2}F]_{C^{\beta}(B_{d_{\Sigma}(x)/4}(x))}\leq\tilde{C}\|\xi\|_{L^{\infty}(B_{d_{\Sigma}(x)/2}(x))},$
for any $x\in\Sigma$, then
(4.6) $d_{\Sigma}|\nabla_{\theta}h|+d^{2}_{\Sigma}|\nabla^{2}_{\theta}h|+d_{\Sigma}^{2+\beta}[\nabla^{2}_{\theta}h]_{C^{\beta}(B_{d_{\Sigma}(\cdot)/2}(\cdot))}\leq C\xi\quad\text{on }\Sigma,$
where $C$ is a positive constant depending only on $n$, $k$, $\beta$, $\Sigma$, $\bar{C}$, and $\tilde{C}$.
@@
Set $\rho_{0}=\sqrt{n/(2k^{2})}$. Note that on $\{\rho<\rho_{0}\}$,
(4.7) $-\frac{n(n+2)}{4}+\Big(k-\frac{n}{2}+1\Big)\Big(k+\frac{n}{2}-1\Big)\rho^{2}+\frac{(n-2)^{2}}{4}\rho^{2}+\frac{n(n-1)}{4}$
$\qquad=-\frac{3n}{4}+k^{2}\rho^{2}<-\frac{3n}{4}+\frac{n}{2}=-\frac{n}{4}<0,$
@@
(4.8) $\begin{cases}\begin{aligned} c&=-\frac{n(n+2)}{4}+\sigma_{k}\rho^{2}\quad\text{on }\{\rho<\rho_{0}\},\\ c&\leq-\frac{(n-2)^{2}}{4}\rho^{2}-\frac{n(n-1)}{4}\quad\text{on }\Sigma.\end{aligned}\end{cases}$
Step 1. Set $L_{0}=\rho^{2}\Delta_{\theta}+c$, we firstly solve
(4.9) $L_{0}\tilde{h}=\rho^{2}F.$
@@
Now, we can take $\psi=\xi$ and $\delta=n/4$ in Theorem A.1. Hence, we obtain an $\tilde{h}\in C^{\infty}(\Sigma)$ satisfying (4.9) with
(4.10) $\Big\|\frac{\tilde{h}}{\xi}\Big\|_{L^{\infty}(\Sigma)}\leq\frac{1}{\delta}\Big\|\frac{\rho^{2}F}{\xi}\Big\|_{L^{\infty}(\Sigma)}\leq C,$
@@
Step 2. Set $L_{1}=\Delta_{\theta}-\frac{n(n+2)}{4\rho^{2}}$, we secondly solve
(4.11) $L_{1}\bar{h}+\sigma_{k}\bar{h}=G,$
@@
$G=\sigma\phi_{1}+\tilde{G}$,
with
$\tilde{G}=\tilde{h}\Big(\frac{c}{\rho^{2}}+\frac{n(n+2)}{4\rho^{2}}-\sigma_{k}\Big).$
By (4.8), we have $\tilde{G}=0$ on $\{\rho<\rho_{0}\}$ and $\tilde{G}\in C^{\infty}(\bar{\Sigma})$.
(i) If $\sigma_{k}<\lambda_{1}$, we take $\sigma=0$.
(ii) If $\sigma_{k}=\lambda_{1}$, we take $\sigma=-(\tilde{G},\phi_{1})_{L^{2}(\Sigma)}$, then $(G,\phi_{1})_{L^{2}(\Sigma)}=0$.
Hence, in both cases, we can apply Lemma A.2(2) with $\lambda=\sigma_{k}$ and $f=G$ to find a $\bar{h}\in C^{\infty}(\Sigma)$ satisfying (4.11) with
$\|\bar{h}\|_{H^{1}_{0}(\Sigma)}\leq C\|G\|_{L^{2}(\Sigma)}\leq C\|\tilde{G}\|_{L^{2}(\Sigma)}\leq C.$
@@
(4.12) $\|\bar{h}\|_{L^{\infty}(\Sigma)}\leq C(\|\bar{h}\|_{L^{2}(\Sigma)}+1)\leq C.$
Step 3. Take $h=\tilde{h}+\bar{h}$. By (4.9) and (4.11), we can verify that $h\in C^{\infty}(\Sigma)$ satisfies (4.1). In addition, by combining (4.10) and (4.12), we obtain (4.4). Next, we assume that $F$ satisfies (4.5). Recall $\xi=\rho^{-(n-2)/2}\sim d_{\Sigma}^{-(n-2)/2}$, and we note for any $z\in B_{d_{\Sigma}(x)/2}(x)$,
@@
By applying the scaled $C^{2,\alpha}$-estimate to (4.1) and using (4.3)-(4.5) and (A.12), we get
${d_{\Sigma}(x)\|\nabla_{\theta}h\|_{L^{\infty}(B_{d_{\Sigma}(x)/8}(x))}+d^{2}_{\Sigma}(x)\|\nabla^{2}_{\theta}h\|_{L^{\infty}(B_{d_{\Sigma}(x)/8}(x))}+d_{\Sigma}^{2+\beta}(x)[\nabla^{2}_{\theta}h]_{C^{\beta}(B_{d_{\Sigma}(x)/8}(x))}}$
$\qquad\leq C(\|h\|_{L^{\infty}(B_{d_{\Sigma}(x)/4}(x))}+\frac{d^{2}_{\Sigma}(x)}{d^{2}_{\Sigma}(x)}\|\rho^{2}(F+\sigma\phi_{1})\|_{L^{\infty}(B_{d_{\Sigma}(x)/4}(x))}$
$\qquad\qquad+\frac{d^{2+\beta}_{\Sigma}(x)}{d^{2}_{\Sigma}(x)}[\rho^{2}(F+\sigma\phi_{1})]_{C^{\beta}(B_{d_{\Sigma}(x)/4}(x))})$
@@
Remark 4.2.
In particular, when $k=1$, we have $\sigma_{1}=-n(n-4)/4$. Note that $n(n-4)/4\leq 0<\lambda_{1}$ for $n\geq 4$ and $n(n-4)/4=3/4<\lambda_{1}$ for $n=3$ by Lemma A.2(1) and Lemma A.2(3).
We further observe that when $\sigma_{k}=\lambda_{1}$ for some integer $k>1$, the solution of equation (4.1) is not unique. In fact, if $h$ is a solution of (4.1), then $h+t\phi_{1}$ is also a solution for any $t\in\mathbb{R}$.
@@
Let $\lambda_{1}$ be the first eigenvalue of $L_{1}$ in (A.8) and let $\phi_{1}\in H^{1}_{0}(\Sigma)\cap C^{\infty}(\Sigma)\cap C(\bar{\Sigma})$ be the corresponding eigenfunction satisfying $\phi_{1}>0$ in $\Sigma$ with $\|\phi_{1}\|_{L^{2}(\Sigma)}=1$. Note that the existence of such a function $\phi_{1}$ is guaranteed by Lemma A.2, and $|\phi_{1}|\leq C^{\prime}\xi$ for some constant $C^{\prime}>0$ depending only on $n$, $\lambda_{1}$, and $\Sigma$.